What concentration of HOCI is required to oxidize 20mg/L NH3-N to N2?


1
Expert's answer
2021-11-08T08:22:02-0500


Molar mass of NH3= 17.031

Molar mass of HOCl=52.46 g/mol

Molar mass of N2= 28.0134 g/mol

= 20/17.031

= 1.174

= 52/28.0134= 1.86

= 1.174×1.86= 2.179mol/L